OpenStudy (anonymous):

what is the solution to f(x)=2.4+0.4x g(x)=0.20x-1.2

OpenStudy (shiraz14):

Is your question for f(x)=g(x)?

OpenStudy (anonymous):

yes

OpenStudy (shiraz14):

2.4+0.4x=0.2x-1.2 0.2x=-3.6 x=-18